OpenStudy (anonymous):

a six sided die is rolled 24 times If the die was rolled 96 more times and the pattern continues, what is the range representing the frequency 6 will most likely be rolled in the 120 total rolls?

OpenStudy (anonymous):

Side # of Times Landed On 1 2 2 6 3 2 4 6 5 5 6 3

OpenStudy (anonymous):

i believe the answer is between two numbers

OpenStudy (anonymous):

a few ways to do this. If you just take 3 times 5, you get 15

OpenStudy (anonymous):

since 24*5 = 120

OpenStudy (anonymous):

the most proper way to go about it is to take the proportion you've been getting...in this case you get a 6 every 1 out of 8 times

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

1/8 • 120 = x/120 • 120 x = 15 so between 13 and 17 thanks

OpenStudy (anonymous):

15 is correct as the most likely number, but out of curiosity, how did you get that range? Were you choosing from a list of options? Or is there some mathematical reason that you say that range?

OpenStudy (anonymous):

yeah because it was actually asking me between what numbers would the actual answer be

OpenStudy (anonymous):

... Was this multiple choice?

OpenStudy (anonymous):

Why do you say 13 and 17 rather than 14 and 16 or 12 and 18 or 11 and 19?

OpenStudy (anonymous):

because it explained it would be by pairs lik2,4,6,8 etc in a pair of three numbers the middle being the answer it explained